find the equation in slope intercept form: slope: -1 ordered pair (2,4)
Answer by mananth(16949)

The equation in point slope form is
y-y1=m(x-x1)
m=-1
(x1,y1)=> (2,4)
y-4 = -1(x-2)
y-4=-x+2
y=-x+2+4
y=-x+6